A dealer sells an article for GH¢600. He will allow you to buy it for GH¢240 down, with

the balance to be paid in GH¢30 monthly installments for a year. If you pay cash, he will

give you a 10% discount.

a. what is the interest rate j12 paid by the purchaser who uses the installments plan?

b. what is the annual effective rate being charged on the installment plan?



1
Expert's answer
2021-04-26T06:04:51-0400

Selling price=600

Pay cash discount= 10%

Instalment purchase

Deposit=240

30 monthly installments=360/30=12 per installment.

a.) Interest rate on installment is the cost forgone by not paying on cash

Interest = 10%

b.) Annual effective interest rate

"Annual=(1+(\frac{nominal rate}{no of periods}))^{no of periods} -1"

"Annual=(1+(\frac{0.1}{30}))^{30} -1=0.104976133"

"Annual rate=0.104976133\times100=10.498"